University of Colorado at Boulder University of Colorado CU Home Search A to Z Index Map
Laboratory for Atmospheric and Space Physics

X-2 Flare

1.0 Summary of Significant Activities for the Week

POLAR spacecraft operations were not routine for this reporting period.

On May 6 at 08:15:27 UT, the POLAR spacecraft was upset by a solar proton
event associated with an X-2 class flare. On data acquisition at 10:00 UT,
all spacecraft and instrument telemetry were found to be unintelligible. An
unscheduled DSN contact was immediately arranged to assess the status of
spacecraft subsystems and instruments. The ability of the spacecraft to
accept commands was verified by executing command macros stored on the
spacecraft. Verification of command receipt and execution was verified by
noting the presence or absence of the garbled telemetry.

A reset of the spacecraft telemetry module (GTM) was accomplished by
reloading the module's code from ROM. On recovery of valid telemetry, the
magnetometer (MFE) was found to have suffered a single event upset (SEU).
Invalid magnetics data being sent to the spacecraft processor (CAP) for
computation of magnetics azimuth and elevation (Magaz/Magel) caused an
overflow error in the CAP. This error was cleared by turning off the flow
of the high rate data from MFE to the CAP. In addition, because of the
interruption of telemetry caused by resetting the GTM, the HYDRA instrument
automatically entered a safe mode (high and low voltage off, main power
remains on). SEPS also entered a safe mode. The TIDE and TIMAS instruments
recorded the interrupt in telemetry as queuing errors, with no affect on
the instruments.

The spacecraft and instruments were subsequently returned to normal
operations at the scheduled DSN contact at 15:40 UT. However, anomalous
temperature readings persist in the housekeeping telemetry for HYDRA,
CEPPAD and MFE, suggesting that a telemetry multiplexer was also upset. The
specific multiplexer has not yet been identified. Correct readings of the
temperatures in question for MFE and CEPPAD (a single parameter in each
case) appear in each instrument's telemetry allocation, as distinct from
the spurious values occurring in the spacecraft housekeeping telemetry
allocation. Anomalous HYDRA temperatures remain in both instrument and
spacecraft telemetry allocations.

The Plasma Source Instrument (PSI), part of the TIDE experiment, could not
be ignited as planned May 6. A second attempt at ignition on May 7 was
aborted during the initial bakeout cycle due to an anomalous hight voltage
reset of the TIDE instrument. TIDE was returned to full operation on May 8.
Another attempt at PSI ignition will occur on May 11.

The TIDE instrument suffered a high voltage reset on May 7 from an
unidentified source. The recovery to full operation was executed on May 9.
In association with the TIDE reset, the HYDRA eeprom write-protect
circuitry was enabled. In addition, the VIS heater current on the pulse bus
showed a dropout. The source of the anomaly is under investigation.

Data capture - Week 17 (DOY 117-123): 100.0% Year to date : 99.8%.

2.0 Upcoming Events

The last eclipse of this season will occur on March 13.

3.0 Instrument Status

On a routine daily basis, instruments execute orbit position-dependent
commands from the stored table. Commanding for the platform instruments is
dictated by the daily platform pointing plan. Non-routine activities are
indicated below.

HIT             Fully Operational
MICS            Fully Operational
IPS             Fully Operational
IES             Fully Operational
HIST            Fully Operational
Events/Comments: Reconfiguration not required for maneuvers.
EFI             Fully Operational

HYDRA           Fully Operational
Events/Comments: EEPROM write-protect enable and error clear.

MFE             Fully Operational
Events/Comments Instrument recovery following single event upset on May 6.

PIXIE Anomalous power supply behavior in front detector limits that
detector's operating time.
Events/Comments: Duty cycling of the front detector from the stored command
PSI             Fully Operational. Part of TIDE experiment.
Events/Comment: Unsuccessful ignition attempt on May 6. Second attempt on
                May 7 aborted due to anomalous TIDE HV reset.

PWI             Non-operational

SEPS            Fully Operational
Events/Comments: Processor reset following single event upset on May 6.

TIDE            Fully Operational
Events/Comments: Full HV recovery following nomalous reset on May 7
Reset telemetry queuing error flag.

TIMAS           Fully Operational
Events/Comments: Reset telemetry queuing error flag.

UVI             Fully Operational on B-side.
Events/Comments: .

VIS             Fully Operational

4.0 Spacecraft Status

Spacecraft status is detailed in Attachment A.

5.0 Problem Reports For This Week

Eleven problem reports were written during the reporting period. A summary
listing of the reports is provided as Attachment B.

1. April May 4. Station 46 was unable to provide 3 kw uplink or range due
to inoperable PA1 and PA2. The PA bypass (16 watts) was used to overcome
the problem.

2. May 6. All spacecraft and instrument telemetry invalid at AOS. Suspected
SEU of telemetry subsystem (GTM). Per established contingency procedure,
GTM reset performed in the blind. Valid telemetry restored. Through
subsequent analysis of tape dump, SEU determined to occur at 126:08:15:27.

3. May 7. Degraded data on Channel 1from 00:30 - 01:00 due to TGC-5 problem
at Station 66. Station switched to TGC-3 to recover.

4. May 7. A TIDE high voltage telemetry indicated a reset condition at
22:13:22. This occurred during bakeout of the Plasma Source Instrument
associated with the TIDE experiment. PSI powered off. TIDE HV recovery
scheduled for later contact.

5. May 7. HYDRA eeprom write-protect flag "Enabled" at same time as TIDE
reset. Flag reset to "Disabled" under direction of instrumenter.

6. May 8. Flight software single bit error noted at AOS (128:07:10). Error
cleared using standard procedure.

7. May 8. Flight software single bit error noted at AOS (128:15:50). Error
cleared using standard procedure.

8. May 8. Real-time and playback data lost from 16:20:00to 16:22:00 due to
loss of contact by Station 16. Station performed a re-sweep to recover
contact. FOT repositioned tape recorder before resuming playback. Data are

9. May 8. Flight software single bit error noted at 128: 01:55:22. Error
cleared using standard procedure.

10 May 8. Flight software single bit errors noted at 129:07:10 and
129:01:20. Errors cleared using standard procedure. SEPS real-time
commanding interrupted.

11. May 10. Station 66 had problems establishing 2-way configuration. Noted
that spacecraft was in non-coherent mode. Executed on-board macro to place
spacecraft in coherent mode. Station acquired spacecraft to proceed with
normal contact.